# ConsDreamer: Advancing Multi-View Consistency for Zero-Shot Text-to-3D Generation

### v1

@@ IV-C Application of the VDM to T2I Tasks
2) Quantitative Results: We compared the success rate of generating images with specific views using Stable Diffusion and our method. Given the open-ended nature of the generated content, a strict definition was adopted for counting success cases. Images exhibiting hallucinations (e.g., artifacts or incorrect semantics) were excluded. Additionally, in instances where multiple entities appeared in an image, it was considered a failure if even one entity did not match the specified view. The results, presented in Table III, demonstrate that our method achieves a significantly higher success rate compared to Stable Diffusion, underscoring its effectiveness in aligning multimodal view semantics.
TABLE III: Success Rate of Specific View Generation
Methods Successful Generation Rate(%$\uparrow$
@@
Stable Diffusion [34] 66.3 28.7
Stable Diffusion+VDM 80.2 77.2
  • $\uparrow$
    Higher scores indicate better performance.
@@ V Conclusion
In this paper, we conducted a mathematical analysis of the multi-face Janus problem in zero-shot text-to-3D generation. To address this challenge, we proposed a novel framework, $ConsDreamer$, which incorporates a View Disentanglement Module to effectively eliminate prior view biases and enhance view control. Additionally, we designed a Partial Order Loss $\mathcal{L}_{P}$ to explicitly enforce multi-view consistency across generated content. We validated the significant improvements of our method over SOTA approaches through comprehensive qualitative and quantitative evaluations. Beyond its application in 3D generation, the View Disentanglement Module also proves effective in 2D generation tasks, correcting view-related semantic biases during the image synthesis process. These contributions collectively establish a robust framework for generating 3D content with greater consistency and realism, paving the way for broader applications in creative industries, virtual reality, and beyond.

### v2

@@ IV-C Application of the VDM to T2I Tasks
2) Quantitative Results: We compared the success rate of generating images with specific views using Stable Diffusion, Perp-Neg and our method. Given the open-ended nature of the generated content, a strict definition was adopted for counting success cases. Images exhibiting hallucinations (e.g., artefacts or incorrect semantics) were excluded. Additionally, in instances where multiple entities appeared in an image, it was considered a failure if even one entity did not match the specified view. The results, presented in Table IV, show that our proposed VDM improves rear-view generation success by approximately 36.7% over Perp-Neg, demonstrating that our method achieves a substantially higher success rate and effectively aligns multimodal view semantics.
Figure 14: Generation failure case examples. For complex or multi-object prompts (top), the underlying optimisation may still produce distorted or entangled geometry, while for single-object cases (middle and bottom), ConsDreamer improves view consistency but cannot fully remove illumination artefacts such as oversaturation and colour shifts.
TABLE IV: Success Rate of Specific View Generation
[…]
@@
Stable Diffusion 66.3 28.7
Stable Diffusion + Perp-Neg 77.4 56.5
Stable Diffusion + VDM (Ours) 80.2 77.2
IV-D Limitations
Methods built on score-distillation pipelines, while capable of substantially improving multi-view consistency, still inherit the fundamental limitations of this optimisation paradigm. For highly complex text descriptions or cluttered multi-object scenes, the underlying optimisation may converge to sub-optimal 3D configurations, leading to fragmented or entangled structures. In these regimes, partial-order constraints act only as a relatively weak regularisation signal and cannot by themselves guarantee globally coherent geometry, as illustrated in Fig. 14. Moreover, score-distillation-based approaches predominantly target view-level semantics and geometry, and typically do not explicitly account for illumination artefacts introduced during distillation. Brightness accumulation and colour saturation anomalies can therefore persist even when view consistency is improved (Fig. 14). A principled treatment of such photometric effects would likely require revisiting the distillation objective itself, which we regard as an important direction for future work.
@@ V Conclusion
In this paper, we conducted a mathematical analysis of the multi-face Janus problem in zero-shot text-to-3D generation. While Previous methods alleviate this problem by orthogonalizing negative prompts in the denoising score space, they only operate on the UNet output and overlook the aggregated prior view features formed during the UNet processing, leading to the recurrence of inconsistency when lifting 2D representations to 3D. To address this fundamental limitation, we propose ConsDreamer, which offers two key advantages: (1) a View Disentanglement Module that eliminates view biases in conditional prompts by decoupling irrelevant view components and injecting precise view control prior to denoising; and (2) a similarity-based partial order loss that enforces geometric consistency in the unconditional term by aligning cosine similarities with azimuthal relationships, synergistically enhancing view semantic clarity. Demonstrating strong versatility, ConsDreamer can be seamlessly integrated into diverse 3D representations (e.g., NeRF, 3DGS) and various score distillation paradigms. Extensive experiments across multiple SOTA baselines confirm that our method effectively mitigates the multi-face Janus problem while outperforming existing approaches in both visual quality and multi-view consistency.